# Schema Augmentation for Zero-Shot Domain Adaptation in Dialogue State Tracking

### v1

@@ 8 Conclusion
We developed Schema Augmentation, a data augmentation technique for zero-shot domain adaptation in dialogue state tracking. To assess its effectiveness, we introduced Target Goal Accuracy (TGA), a metric that evaluates performance specifically on unseen target domains. Our method significantly boosts TGA compared to baseline approaches across two widely-used datasets, demonstrating robustness to out-of-distribution information in prompts during test time. Ablation studies further illuminate the key factors contributing to these improvements. These results underscore the potential of Schema Augmentation as a simple yet powerful tool for enhancing domain adaptation in task-oriented dialogue systems. We recommend future work to explore similar augmentation strategies to improve generalization to unseen domains in other applications.
@@ Appendix A Appendix
A.1 Hyperparameters
@@
A.2 Compute
@@
A.3 Slots and Domains
Tables 2 and 3 show the domain and slot combinations for the two datasets. Taxi, train, and bus were chosen as holdout domains due to many slots in common with each other and few slots in common with other domains.
attraction bus hospital hotel police restaurant taxi train
address
@@ A.4 Replacements
A.5 Schema Augmentation Examples
@@
A.6 Licenses
All datasets and models used are free and open-source. MultwiWOZ and SpokenWOZ are available under MIT license. Gemma-2 and Mistral are available under the Apache 2.0 license. All models and datasets are intended for reseach purposes, which is consistent with this work. Datasets are widely used and reported not to contain personal information or offensive content.

### v2

@@ 6 Conclusion
We developed Schema Augmentation, a data augmentation technique for zero-shot domain adaptation in dialogue state tracking. To assess its effectiveness, we introduced Target Goal Accuracy (TGA), a metric that evaluates performance specifically on unseen target domains. Our method significantly boosts TGA compared to baseline approaches across two widely-used datasets, demonstrating robustness to out-of-distribution information in prompts at inference.
@@ Appendix A Appendix
A.1 Ablation Study
Baseline(no aug.)SSAESA1010101020202020303030304040404018.218.218.218.234.534.534.534.540.740.740.740.710.410.410.410.42020202016.516.516.516.5TGA (%) vs. ExperimentOriginalShuffled Descriptions/Values
Figure 2: Ablation results for MultiWOZ 2.2. SSA and ESA are our Schema Augmentation methods. No Aug means fine-tuning gemma-2-9b-it without data augmentation, to illustrate the effects of Schema Augmentation.
We conduct an ablation study to shed light on the mechanism of Schema Augmentation, and why the encoding variant generally outperforms the synonym variant. In our ablation, we introduce a mismatch between slots and their corresponding descriptions/values in the schema by random shuffling, i.e. every slot is paired with a randomly chosen description and value list from a different slot during preprocessing. The shuffling is randomized on every sample to ensure there is no correlation between description/values and the correct slot in the answer. We conduct our study on the MultiWOZ 2.2 dataset, which was preferred over 2.1 due to its higher quality and lower noise content. Figure 2 shows the ablation results. We compared our methods SSA and ESA to a baseline of standard fine-tuning without data augmentation, but using the same hyperparameters and model (gemma-2-9b-it). Upon shuffling descriptions/values, we observe the greatest drop in TGA for ESA (-24.5%), and second greatest for SSA (-14.5%). The baseline had the lowest drop (-7.8%). A higher drop in TGA means a greater desgradation in performance when the information content of descriptions/values is effectively removed. This suggests that Schema Augmentation encourages the model to pay more attention to the descriptions and values of the slots, and that effect is stronger with ESA than SSA.
A.2 Computing TGA
Prior work uses the "leave-one-out" cross-domain adaptation setup Lin et al. (2021a, b). In this scenario, a single domain at a time is designated as the holdout domain $d_{H}$ and withheld from the training set. A model is trained on the remaining four domains, and then tested on the test split. Prior work reports JGA for $d_{H}$, computed over dialogues that include $d_{H}$. We will refer to this as holdout JGA, or $\text{JGA}^{H}$.
In our experiments, we designated taxi and train as holdout domains due to slot overlap (see Section 4.1). In order to convert $\text{JGA}^{H}$ to TGA, we will combine $\text{JGA}^{\texttt{taxi}}$ and $\text{JGA}^{\texttt{train}}$ using the known test set distribution over domains:
$\text{JGA}^{\texttt{taxi}+\texttt{train}}=\text{JGA}^{\texttt{taxi}}\cdot N_{\texttt{taxi}}$
$+\text{JGA}^{\texttt{train}}\cdot N_{\texttt{train}}$
We also need to recognize that unlike TGA, $\text{JGA}^{H}$ includes empty states. This is due to the fact that the domain filtering is done per dialogue, but the data itself is per turn. For example, a dialogue labelled with the taxi domain can have many turns which have empty taxi slots - all the turns which appear before taxi is metioned in the dialogue. TGA is a stricter metric that does not consider these empty states. In order to convert to TGA, we thus need an additional assumption about the performance on empty states. In all experiments done in this work, the accuracy of empty states was >99% on all domains for both datasets. Thus, without a significant loss in accuracy, we assume a 100% accuracy of prior work on empty states. This will not be perfectly accurate, but without generations provided by prior work, we cannot perfectly determine this accuracy. With this assumption, we can now estimate TGA:
$\text{TGA}^{\texttt{taxi},\text{est}}=\frac{\text{JGA}^{\texttt{taxi}}\cdot N_{\texttt{taxi}}-N_{\texttt{taxi}}^{\text{empty}}}{N_{\texttt{taxi}}^{\text{nonempty}}}$ (8)
The formula for $\text{TGA}^{\texttt{train},est}$ is the same, and we combine them to get our overall estimate of TGA:
$\text{TGA}^{\text{est}}$ $=\text{TGA}^{\texttt{taxi},\text{est}}\cdot N_{\texttt{taxi}}^{\text{nonempty}}$
$+\text{TGA}^{\texttt{train},\text{est}}\cdot N_{\texttt{train}}^{\text{nonempty}}$ (9)
A.3 Dataset Description
MultiWOZ is a multi-domain task-oriented dialogue dataset comprising annotated dialogues across eight domains including hotel booking, restaurant reservation, and taxi ordering. Each dialogue is annotated with the dialogue state at each turn. For consistency with prior work, we run our main experiments on MultiWOZ 2.1 Eric et al. (2019), a popular revision of the original. We also report results on SpokenWOZ Si et al. (2024). SpokenWOZ dialogues were collected from crowdworkers engaging in spoken conversations and includes text transcriptions from an automatic speech recognition (ASR) system. We perform our experiments on the audio transcriptions. For ablation studies, we use a cleaner version of MultiWOZ, MultiWOZ 2.2 Zang et al. (2020) due to the lower noise content.
A.4 Hyperparameters
@@
A.5 Mistral Results
Our results on the Mistral 7B model from Table 4 corroborate our findings that the proposed augmentation methods yield significant gains for dialog state tracking when applied to multiple different LLMs.
Method $\textbf{JGA}_{\text{taxi}}$ $\textbf{JGA}_{\text{train}}$ TGA
MultiWOZ 2.2
SSA 65.0 42.9 24.9
ESA 67.9 42.6 26.0
SpokenWOZ
SSA 66.7 35.2 12.5
ESA 66.7 37.7 15.4
Table 4: MultiWOZ 2.2 and SpokenWOZ results using the Huggingface implementation of the Mistral-7B-Instruct-v0.3 model.
A.6 Compute
@@
A.7 Slots and Domains
Tables 5 and 6 show the domain and slot combinations for the two datasets. Taxi, train, and bus were chosen as holdout domains due to many slots in common with each other and few slots in common with other domains.
attraction hotel restaurant taxi train
@@ A.8 Replacements
A.9 Schema Augmentation Examples
@@
Listing 6: Prompt Example
### Instructions: Give the dialogue state at the end of the given dialogue, formatted in JSON. Follow the schema and only use the given pre-defined slots and their possible values. ‘Possible values: []‘ means open-ended (the slot can take on any value). Omit any slots with empty values from your answer. If no slots can be filled from the dialogue, respond with an empty json object.
### Schema:
- Slot: attraction-area; Description: area to search for attractions; Possible values: [’centre’, ’east’, ’north’, ’south’, ’west’]
- Slot: attraction-name; Description: name of the attraction; Possible values: []
- Slot: attraction-type; Description: type of the attraction; Possible values: [’architecture’, ’boat’, ’cinema’, ’college’, ’concerthall’, ’entertainment’, ’museum’, ’multiple sports’, ’nightclub’, ’park’, ’swimmingpool’, ’theatre’]
- Slot: attraction-entrancefee; Description: how much is the entrance fee; Possible values: []
- Slot: attraction-openhours; Description: open hours of the attraction; Possible values: []
- Slot: attraction-address; Description: address of the attraction; Possible values: []
- Slot: attraction-phone; Description: phone number of the attraction; Possible values: []
- Slot: attraction-postcode; Description: postal code of the attraction; Possible values: []
### Dialogue:
USER: i am looking for a college type attraction .
SYSTEM: there are 18 colleges i have found , would you prefer 1 in town centre or in the west ?
USER: i would like to visit on in town centre please .
A.10 Licenses
MultiWoz is available under an MIT License, and SpokenWoz is available under a CC-BY-NC 4.0 license. Gemma-2 is available pursuant to the Gemma Terms of Use. Mistral is available under an Apache 2.0 license. All models and datasets are intended for research purposes, which is consistent with this work.
Datasets are widely used and reported not to contain personal information or offensive content.